# Choose Wisely: Data-Enabled Predictive Control for Nonlinear Systems Using Online Data Selection

### v1

@@ Abstract
This paper proposes Select-Data-Enabled Predictive Control (Select-DeePC), a new method for controlling nonlinear systems using output-feedback for which data are available but an explicit model is not. At each timestep, Select-DeePC employs only the most relevant data to implicitly linearize the dynamics in “trajectory space.” Then, taking user-defined output constraints into account, it makes control decisions using a convex optimization. This optimal control is applied in a receding-horizon manner. As the online data-selection is the core of Select-DeePC, we propose and verify both norm-based and manifold-embedding-based selection methods. We evaluate Select-DeePC on three benchmark nonlinear system simulators—rocket-landing, a robotic arm and cart-pole inverted pendulum swing-up—comparing them with standard DeePC and Time-Windowed DeePC, and find that Select-DeePC outperforms both methods. The source code can be found at: https://github.com/naefjo/choose-wisely-paper
Machine Learning, ICML
@@ 1 Introduction
Model-based methods have established their value in the Reinforcement Learning (RL) community, even for highly nonlinear systems (Moerland et al., 2022). From a different approach, the Control community has developed interest in “Direct” Data-Driven Predictive Control methods (Dörfler et al., 2023), which typically make Linear, Time-Invariant (LTI) assumptions. This paper proposes Select-Data-enabled Predictive Control (Select-DeePC), a Direct Data-Driven Predictive Control method for nonlinear systems.
Output or safety constraints are challenging for standard RL methods, which learn the constraints by exhaustive trial and error (Sutton & Barto, 1998). Output-constrained systems have traditionally been the domain of Model Predictive Control (MPC) (Morari & H. Lee, 1999), which solves a receding-horizon optimal control problem based on an externally-provided model with an online optimization. In contrast, RL has favored planning methods that are unconstrained and “gradient-free” and, instead of requiring an externally-provided model, use state-action queries of the system. Recently, planning using forward passes through generative or world models have been used to reduce RL sample complexity requirements (Moerland et al., 2022). However, these zero-order-type planning methods usually have no structured way of incorporating output constraints and require the user to offload this complexity into other parts of the method, such as the model itself or the cost.
@@
Figure 1: Closed-loop position trajectories of Select-DeePC (our method), standard DeePC and Time-Windowed DeePC in the MuJoCo Reacher (left) and the Rocket Lander (right) Gym environments, as well as the corresponding open-loop predictions (dashed). In both environments, Select-DeePC converges to the provided setpoint while standard and Time-Windowed DeePC fail to track the setpoints.
@@
Data-Enabled Predictive Control (DeePC) (Coulson et al., 2019) constructs feasible trajectories of an LTI system by linearly combining input-output trajectories in the training data set. DeePC and its variants have demonstrated impressive results on a number of control problems. However, the LTI restriction is severe.
@@
Select-DeePC, proposed in this paper, extends DeePC to the nonlinear setting by iteratively optimizing a convex quadratic program over an implicit linearization in “trajectory space” constructed from data. We define trajectory space as the high-dimensional space in which each dimension corresponds to an input or output over the course of a discrete-time, finite-length trajectory. Linearizing in trajectory space is achieved by selecting only trajectories that are close to the current operating point in trajectory space. Thus, Select-DeePC is a data-driven alternative to nonlinear MPC solved via Sequential Quadratic Programming (SQP), which calculates a trajectory space linearization at each timestep from an externally provided model. SQP-MPC is a standard model-based optimal control tool, commonly used in both academia and industry (Verschueren et al., 2020).
Literature Review DeePC was first introduced by Coulson et al. as an alternative to the sequential, indirect system identification and MPC design pipeline (2019). It uses Willems’ fundamental Lemma (2005) for deterministic LTI systems to generate trajectories of the system as opposed to an identified explicit linear model and is hence considered a direct method. Such predictions fits into the behavioral control framework, see, for instance, (Willems, 1991). For a survey of DeePC we refer the interested reader to (Markovsky et al., 2023).
Several extensions to DeePC have been proposed to handle nonlinear systems in a structured way as opposed to treating the observed nonlinearity as measurement noise on a linear system. Berberich and Allgöwer provide a list of tailored formulations of the Fundamental Lemma to specific classes of nonlinear systems (2024). These formulations enable efficient construction of a DeePC formulation for systems belonging to these specific system classes. However, this already imposes a strong prior on the global nonlinear structure of the dynamical process, which might not be available for an arbitrary system. Alternatively, Berberich et al. propose using a set of the most recent past observations along the current closed-loop trajectory to form the basis of a linear predictor (2022). This predictor is continually updated as new data is obtained in a sliding window fashion, effectively forming a locally linear approximation of the underlying system. As this method only uses recent closed-loop data, it is practical only for systems with benign nonlinearities and with very high signal-to-noise ratios.
Other approaches to nonlinear-DeePC include reformulating the predictor in a kernelized fashion, enabling the use of nonlinear kernels (Lian & Jones, 2021; Huang et al., 2024), using nonlinear basis functions or reformulating the problem in a lifted state-space using Koopman operators (Lazar, 2023). While these methods enjoy greater flexibility due to their use of nonlinear function approximators, the resulting optimization becomes high-dimensional, nonlinear, and generally nonconvex, requiring increased online computational complexity and resulting in suboptimal outcomes.
@@
Figure 2: Venn-diagram demonstrating how the discussed controllers relate and their strengths and weaknesses. Select-DeePC combines the best of all three worlds by bridging the gap of explicit constraint handling without the requirement for Jacobians of the model achieved by direct methods.
@@
The premise of DeePC, solving an optimal control problem based purely on samples of a generative model without requiring gradient information, fits well into the landscape of planning algorithms in model-based RL. Optimization-based planning techniques suffer from accumulating prediction errors (Moerland et al., 2022), exploding gradients due to backpropagation through time or model exploitation in regions of low data density (Kurutach et al., 2018). Currently predominant planning techniques, herein called Sampling-Based MPC, are gradient-free methods that instead rely on sampling of the forward process model, making them computationally cheap to evaluate. Nagabandi et al. propose to generate rollouts of the forward dynamics from random control inputs (2018). Then, the rollouts are scored according to some reward function and the first action of the best trajectory is applied. A more sophisticated version of this same concept is the cross-entropy method for optimization, which uses iterative refinement of the probability distribution over the control inputs (Botev et al., 2013). A related algorithm is Model Predictive Path Integral Control (MPPI) (Williams et al., 2017), which can be interpreted as an importance sampling scheme, weighing the control actions according to their optimality. Sampling-Based MPC in conjunction with RL methods have been successfully demonstrated in various algorithms and applications such as but not limited to (Chua et al., 2018; Hafner et al., 2019; Nagabandi et al., 2019; Kahn et al., 2020). Other recent approaches include using a generative model to generate trajectories and then using an inverse model to predict the required action (Ajay et al., 2023) or simply rely on random shooting and selecting the action sequence with highest reward (Zhou et al., 2024). Figure 2 shows a schematic overview of the control strategies mentioned and their respective strengths and weaknesses.
@@
While we study the performance of Select-DeePC on a fixed data set, which is common in Data-Driven Control literature, and hereby diverge from the common approach in Sampling-Based MPC where a simulator is queried at each decision moment, we note that Select-DeePC can readily be extended to the simulator-based scenario by, e.g., using random roll-outs from the simulator as the data set.
Contributions. We propose Select-DeePC, a data-driven quadratic program predictive controller for nonlinear systems which implicitly linearizes the system dynamics in the extended input-output space by selecting and using only the most relevant data at each decision moment from a fixed data set. It has the following benefits:
@@
    In contrast with the nonlinear Data-Driven Predictive Control methods in the literature such as basis function methods, Select-DeePC produces trajectories which are consistent with the underlying nonlinear system using a convex quadratic programming solver, yielding a method that can be implemented online in fast dynamic environments.
@@
    In contrast with model-based RL methods, Select-DeePC allows users to explicitly program-in output constraints and change said constraints, as well as cost/reward function, a posteriori. This allows Select-DeePC to be employed in changing safety conditions without requiring retraining or retuning.
@@
    Select-DeePC is easy to implement, requiring only a few lines of code for the end-user. Moreover it is conceptually easy to reason about and adapt due to its modular structure. In particular, it enables a rich choice of data-selection methods.
@@
The core idea behind Select-DeePC is the selection of a subset of data from a large data set at each decision moment. We introduce two methods for the data-selection problem, norm-based and manifold-embedding-based selection, and discuss their relevant strengths and weaknesses.
@@
Finally, we demonstrate the benefits of Select-DeePC on three benchmark nonlinear systems including a planar rocket-landing, a two degree of freedom robotic arm, and a cart-pole inverted pendulum, comparing it to standard DeePC and a Time-Windowed DeePC approach (Berberich et al., 2022) and find that Select-DeePC outperforms both in terms of closed-loop cost and prediction accuracy.
2 Select-DeePC
@@
Select-DeePC is a receding-horizon optimal control approach which determines inputs at each timestep using a predictive optimization. While it does not use a model in the classical sense, in the RL parlance, it could be interpreted as a “model-based” method (Sutton & Barto, 1998), while in the Data-Driven Control community it is a “direct” method (Dörfler et al., 2023).
@@
Select-DeePC derives its name from DeePC which solves a receding-horizon predictive control problem subject to input and output constraints. The novelty of DeePC is that it achieves predictive control purely based on implicit predictions of input-output data without ever constructing a explicit predictor (a model) of future measurements given a set of inputs. This is achieved by linearly combining input-output trajectories $\tau$ of length $L$ from a data set ${\mathcal{D}=\{\tau_{0},\dots,\tau_{n_{d}}\}}$. The $L$-length trajectories $\tau_{i}$ are partitioned into a past trajectory $\tau_{\text{p},i}$ and a future trajectory $\tau_{\text{f},i}$ with lengths $T_{\text{p}}$ and $T_{\text{f}}$ respectively with $T_{\text{p}}+T_{\text{f}}=L$. The implicit predictor in DeePC works by pattern matching the past $T_{\text{p}}$ input-output measurements of the current closed-loop trajectory to a linear combination of the $\tau_{\text{p},i}$ in $\mathcal{D}$. Since this can be achieved by (potentially) many different linear combinations, the objective of DeePC is to then find a linear combination of future trajectories $\tau_{f}$ that minimizes a cost function subject to the pattern matching constraint. Several regularizing terms have been introduced to further bias the choice of said linear combinations. For more details, we refer to Appendix A.
@@
Solving DeePC online using a large data set of trajectories is intractable as the solve time of the optimization problem scales superlinearly with the cardinality of the data set. Furthermore, in the presence of nonlinearities, the LTI assumption of DeePC is violated, resulting in suboptimal performance. We now introduce Select-DeePC, an algorithm that allows us to solve a nonlinear data-driven receding-horizon optimal control problem using a convex quadratic programming routine by choosing, at each solver iteration, a subset of trajectories $\tilde{\mathcal{D}}$ with cardinality $\operatorname{card}(\tilde{\mathcal{D}})=N_{\text{cols}}$ from $\mathcal{D}$ where $N_{\text{cols}}\ll\operatorname{card}\left(\mathcal{D}\right)$.
Algorithm 1 Select-DeePC
1:  function Select-DeePC($u_{\text{p}}$, $y_{\text{p}}$
@@
5:        $u\leftarrow\texttt{DeePC.computeAction}(u_{\text{p}},\ y_{\text{p}},\ \tilde{D})$
@@
Select-DeePC is summarized in Algorithm 1. In between two real-time sampling instances, an iterative procedure is performed where, at every iteration, we select from a data set $\mathcal{D}$ a subset of trajectories $\tilde{\mathcal{D}}$ that are “relevant” to the current linearization point given by the open loop solution $\tilde{\tau}$. While we loosely use the term relevance here, this mechanism will be discussed in more detail in Section 3. Select-DeePC then solves a Data-Driven Predictive Control problem using the subset of data. This is iterated until some convergence criterion is met, e.g. the open-loop prediction $\tilde{\tau}$ converges.
@@
Cost-agnostic. Since the training process/data set generation is structurally independent of the control formulation, a posteriori cost or constraint modification is possible, allowing the method to be applied zero-shot to new problems without the need for retraining.
@@
Hyperarameters. The performance of Select-DeePC depends on the tuning of several hyperparameters. First, the DeePC specific hyperparameters, such as regularization terms and horizon lengths $T_{\text{p}},\ T_{\text{f}}$. These can either be hand-tuned according to established heuristics (Elokda et al., 2021) or optimized through gradient-descent methods (Cummins et al., 2024) or black-box parameter tuning (Berkenkamp et al., 2016). For DeePC tuning guidelines, we refer to (Markovsky et al., 2023). Select-DeePC adds onto these the design parameters $N_{\text{cols}}$ as well as the method of data selection. In the following section, we will investigate two different data-selection methods which subsample a specified number of trajectories from a large offline library of trajectories observed during “training.”
@@
SQP-MPC. The iterative refinement of the solution until convergence is inspired by the model-based SQP-MPC algorithm for nonlinear systems. We direct the interested reader to Appendix B for a detailed discussion of SQP-MPC and the relationship to Select-DeePC. We note that Select-DeePC’s implicit linearization is performed in trajectory space. In the model-based case, linearizing in trajectory space corresponds to a linear time-varying parametrization of the nonlinear dynamics for the MPC. This linear, time-varying parametrization contrasts with an LTI formulation in which the Jacobians are only evaluated at $y(k)$ and kept fixed along the prediction horizon.
@@
Computational complexity. Solving the convex quadratic optimization problem at the basis of Select-DeePC involves repeatedly solving a symmetric system of linear equations (Stellato et al., 2020). Hence, we expect the algorithm to scale with $\mathcal{O}(n^{3})$ where $n$ indicates the number of optimization variables which is directly proportional to the number of trajectories $N_{\text{cols}}$ selected from the full data set. Hence, from a computational complexity standpoint, we wish to perform the optimization with as few trajectories as possible. Additionally, the dominant computation in the data-selection is the sorting of the data according to the relevance to $\tilde{\tau}$ which scales with $\mathcal{O}(n_{d}\log n_{d})$, which we postulate is less significant than the computational complexity of the optimization problem. While Select-DeePC solves a convex quadratic program and is smaller than full-data DeePC or nonlinear DeePC, it still requires more compute than Sampling-Based MPC methods. There exist DeePC variants which reduce the number of optimization variables in the predictive control problem but the comparison of these methods is outside the scope of this study (Favoreel & De Moor, 1999; Breschi et al., 2023). For an overview of existing methods we refer to (Verheijen et al., 2023).

### v2

# Choose Wisely: Data-driven Predictive Control for Nonlinear Systems Using Online Data Selection

@@ Abstract
This paper proposes Select-Data-driven Predictive Control (Select-DPC), a new method for controlling nonlinear systems using output-feedback for which data are available but an explicit model is not. At each timestep, Select-DPC employs only the most relevant data to implicitly linearize the dynamics in “trajectory space”. Then, taking user-defined output constraints into account, it makes control decisions using a convex optimization. This optimal control is applied in a receding-horizon manner. As the online data-selection is the core of Select-DPC, we propose and verify both norm-based and manifold-embedding-based selection methods. We evaluate Select-DPC on three benchmark nonlinear system simulators—rocket-landing, a robotic arm and cart-pole inverted pendulum swing-up—comparing them with standard Data-enabled Predictive Control (DeePC) and Time-Windowed DeePC methods, and find that Select-DPC outperforms both methods.
@@ 1 Introduction
Figure 1: Closed-loop trajectories of Select-DPC (ours), standard DeePC and Time-Windowed DeePC in the MuJoCo Reacher (left) and the Rocket Lander (right) Gym environments, as well as the corresponding open-loop predictions (dashed). In both environments, Select-DPC converges to the provided setpoint while standard and Time-Windowed DeePC fail to track the setpoints.
Model-based methods have established their value in the Reinforcement Learning (RL) community, even for highly nonlinear systems Moerland et al. (2022). From a different approach, the Control community has developed interest in “direct” Data-driven Predictive Control methods Dörfler et al. (2023), which typically make Linear, Time-Invariant (LTI) assumptions. This paper proposes Select-Data-driven Predictive Control (Select-DPC), a direct Data-driven Predictive Control method for nonlinear systems.
Output or safety constraints are challenging for standard RL methods, which learn the constraints by exhaustive trial and error Sutton & Barto (1998). Output-constrained systems have traditionally been the domain of Model Predictive Control (MPC) Morari & H. Lee (1999), which solves a receding-horizon optimal control problem based on an externally-provided model with an online optimization. In contrast, RL has favored planning methods that are unconstrained and “gradient-free” and, instead of requiring an externally-provided model, use state-action queries of the system. Recently, planning using forward passes through generative or world models have been used to reduce RL sample complexity requirements Moerland et al. (2022). However, these zero-order-type planning methods usually have no structured way of incorporating output constraints and require the user to offload this complexity into other parts of the method, such as the model itself or the cost.
@@
The class of Data-driven Predictive Controllers (DPCs) construct feasible trajectories of an LTI system purely from data. Data-enabled Predictive Control (DeePC) Coulson et al. (2019), for instance, linearly combines input-output trajectories in the data set. DPC methods have demonstrated impressive results on a number of control problems. However, the LTI restriction is severe.
@@
Select-DPC, proposed in this paper, extends DPC to the nonlinear setting by iteratively optimizing a convex quadratic program over an implicit linearization in “trajectory space” constructed from data. We define trajectory space as the high-dimensional space in which each dimension corresponds to an input or output over the course of a discrete-time, finite-length trajectory. Linearizing in trajectory space is achieved by selecting only trajectories that are close to the operating point in trajectory space. Thus, Select-DPC is a data-driven alternative to nonlinear MPC solved via Sequential Quadratic Programming (SQP) Diehl et al. (2002), which calculates a trajectory space linearization at each timestep from an externally provided model. SQP-MPC is a standard model-based optimal control tool, commonly used in both academia and industry Verschueren et al. (2020).
@@
Figure 2: Venn-diagram demonstrating the strengths and weaknesses of the discussed methods. Select-DPC combines the best of all three worlds by bridging the gap of explicit constraint handling without the requirement for Jacobians of the model achieved by direct methods.
Literature Review
DeePC was first introduced by Coulson et al. (2019) as an alternative to the sequential, indirect system identification and MPC design pipeline. It uses Willems’ fundamental Lemma Willems et al. (2005) for deterministic LTI systems to generate trajectories of the system as opposed to an identified explicit linear model and is hence considered a direct method. Such predictions fit into the behavioral control framework, see, for instance, Willems (1991). For a survey of DeePC we refer the interested reader to Markovsky et al. (2023).
Several extensions to DeePC have been proposed to handle nonlinear systems in a structured way as opposed to treating the observed nonlinearity as measurement noise on a linear system. Berberich & Allgöwer (2024) provide a list of tailored formulations of the Fundamental Lemma to specific classes of nonlinear systems . These formulations enable efficient construction of a DeePC formulation for systems belonging to these specific system classes. However, this already imposes a strong prior on the global nonlinear structure of the dynamical process, which might not be available for an arbitrary system. Alternatively, Berberich et al. (2022) propose using a set of the most recent past observations along the current closed-loop trajectory to form the basis of a linear predictor, herein referred to as Time-Windowed DeePC. This predictor is continually updated as new data is obtained in a sliding window fashion, effectively forming a locally linear approximation of the underlying system. As this method only uses recent closed-loop data, it is practical only for systems with benign nonlinearities and with very high signal-to-noise ratios.
Other approaches to nonlinear-DeePC include reformulating the predictor in a kernelized fashion, enabling the use of nonlinear kernels Lian & Jones (2021); Huang et al. (2024), using nonlinear basis functions or reformulating the problem in a lifted state-space using Koopman operators Lazar (2023). While these methods enjoy greater flexibility due to their use of nonlinear function approximators, the resulting optimization becomes high-dimensional, nonlinear, and generally nonconvex, requiring increased online computational complexity and resulting in suboptimal outcomes.
@@
The premise of DPC, solving an optimal control problem based purely on samples of a generative model without requiring gradient information, fits well into the landscape of planning algorithms in model-based RL. Optimization-based planning techniques suffer from accumulating prediction errors Moerland et al. (2022), exploding gradients due to backpropagation through time or model exploitation in regions of low data density Kurutach et al. (2018). Currently predominant planning techniques, herein called Sampling-Based MPC, are gradient-free methods that instead rely on sampling of the forward process model, making them computationally cheap to evaluate. Nagabandi et al. (2018) propose to generate rollouts of the forward dynamics from randomly sampled control inputs. Then, the rollouts are scored according to some reward function and the first action of the best trajectory is applied. In a similar vein, the cross-entropy method for optimization uses iterative refinement of the probability distribution over the control inputs to generate an optimal input sequence Botev et al. (2013). A related algorithm is Model Predictive Path Integral Control (MPPI) Williams et al. (2017), which can be interpreted as an importance sampling scheme, weighing the control actions according to their optimality. Sampling-Based MPC in conjunction with RL methods have been successfully demonstrated in various algorithms and applications such as but not limited to Chua et al. (2018); Hafner et al. (2019); Nagabandi et al. (2019); Kahn et al. (2020). Other recent approaches include using a generative model to generate trajectories and then using an inverse model to predict the required action Ajay et al. (2023) or simply rely on random shooting and selecting the action sequence with highest reward Zhou et al. (2024). Figure 2 shows a schematic overview of the control strategies mentioned and their respective strengths and weaknesses.
@@
While we study the performance of Select-DPC on a fixed data set, which is common in Data-driven Control literature, and hereby diverge from the common approach in Sampling-Based MPC where a simulator is queried at each decision moment, we note that Select-DPC can readily be extended to the simulator-based scenario by, e.g., using roll-outs from the simulator as the data set.
Contributions
We propose Select-DPC, a data-driven quadratic program predictive controller for nonlinear systems which implicitly linearizes the system dynamics in the extended input-output space by selecting and using only the most relevant data at each decision moment from a fixed data set. It has the following benefits:
@@
    In contrast with the nonlinear Data-driven Predictive Control methods in the literature such as basis function methods, Select-DPC produces trajectories which are consistent with the underlying nonlinear system using a convex quadratic programming solver, yielding a method that can be implemented online in fast dynamic environments.
@@
    In contrast with model-based RL methods, Select-DPC allows users to explicitly program-in output constraints and change said constraints, as well as cost/reward function, a posteriori. This allows Select-DPC to be employed in changing safety conditions without requiring retraining or retuning.
@@
    Select-DPC is easy to implement, requiring only a few lines of code for the end-user. Moreover it is conceptually easy to reason about and adapt due to its modular structure. In particular, it enables a rich choice of data-selection methods.
@@
The core idea behind Select-DPC is the selection of a subset of data from a large data set at each decision moment. We introduce two methods for the data-selection problem, norm-based and manifold-embedding-based selection, and discuss their relevant strengths and weaknesses. Furthermore, Select-DPC can be used with any DPC method such as Coulson et al. (2019); Breschi et al. (2023); Moffat et al. (2024); Favoreel & De Moor (1999).
@@
Finally, we demonstrate the benefits of Select-DPC on three benchmark nonlinear systems including a planar rocket-landing, a two degree of freedom robotic arm, and a cart-pole inverted pendulum, comparing it to standard DeePC and a Time-Windowed DeePC approach and find that Select-DPC outperforms both in terms of closed-loop cost and prediction accuracy.
2 Select-DPC
@@
Select-DPC is a receding-horizon optimal control approach which determines inputs at each timestep using a predictive optimization. While it does not use a model in the classical sense, in the RL parlance, it could be interpreted as a “model-based” method Sutton & Barto (1998), while in the Data-driven Control community it is a “direct” method Dörfler et al. (2023).
@@
Select-DPC builds upon linear DPC which solves a receding-horizon predictive control problem subject to input and output constraints. The novelty of DPC is that it achieves predictive control purely based on implicit predictions of input-output data without ever constructing an explicit predictor (a model) of future measurements given a set of inputs. DeePC, for instance, achieves this by linearly combining input-output trajectories $\tau$ of length $L$ from a data set ${\mathcal{D}=\{\tau_{0},\dots,\tau_{n_{d}}\}}$. The $L$-length trajectories $\tau_{i}$ are partitioned into a past trajectory $\tau_{\text{p},i}$ and a future trajectory $\tau_{\text{f},i}$ with lengths $T_{\text{p}}$ and $T_{\text{f}}$ respectively with $T_{\text{p}}+T_{\text{f}}=L$. The implicit predictor in DeePC works by pattern matching the past $T_{\text{p}}$ input-output measurements of the current closed-loop trajectory to a linear combination of the $\tau_{\text{p},i}$ in $\mathcal{D}$. Since this can be achieved by (potentially) many different linear combinations, the objective of DeePC is to then find a linear combination of future trajectories $\tau_{f}$ that minimizes a cost function subject to the pattern matching constraint. Several regularizing terms have been introduced to further bias the choice of said linear combinations. For more details, we refer to Appendix A.
@@
Solving DeePC online using a large data set of trajectories is intractable as the solve time of the optimization problem scales superlinearly with the cardinality of the data set. Furthermore, in the presence of nonlinearities, the LTI assumption of DeePC is violated, resulting in suboptimal performance. We now introduce Select-DPC, an algorithm that allows us to solve a nonlinear data-driven receding-horizon optimal control problem using a convex quadratic programming routine by choosing, at each solver iteration, a subset of trajectories $\tilde{\mathcal{D}}$ with cardinality $\operatorname{card}(\tilde{\mathcal{D}})=N_{\text{cols}}$ from $\mathcal{D}$ where $N_{\text{cols}}\ll\operatorname{card}\left(\mathcal{D}\right)$.
Algorithm 1 Select-DPC
1:  function Select-DPC($u_{\text{p}}$, $y_{\text{p}}$
@@
5:        $u\leftarrow\texttt{DPC.computeAction}(u_{\text{p}},\ y_{\text{p}},\ \tilde{D})$
@@
Select-DPC is summarized in Algorithm 1. In between two real-time sampling instances, an iterative procedure is performed where, at every iteration, we select from a data set $\mathcal{D}$ a subset of trajectories $\tilde{\mathcal{D}}$ that are “relevant” to the current linearization point given by the open loop solution $\tilde{\tau}$. While we loosely use the term relevance here, this mechanism will be discussed in more detail in Section 3. Select-DPC then solves a DPC problem using the subset of data. This is iterated until some convergence criterion is met, e.g. the open-loop prediction $\tilde{\tau}$ converges. Notice that there is no restriction on the choice of DPC solver.
@@
Since the training process/data set generation is structurally independent of the control formulation, a posteriori cost or constraint modification is possible, allowing the method to be applied zero-shot to new problems without the need for retraining.
Hyperarameters
@@
The performance of Select-DPC depends on the tuning of several hyperparameters. First, the DPC specific hyperparameters, such as regularization terms and horizon lengths $T_{\text{p}},\ T_{\text{f}}$. These can either be hand-tuned according to established heuristics Elokda et al. (2021) or optimized through gradient-descent methods Cummins et al. (2024) or black-box parameter tuning Berkenkamp et al. (2016). For DPC tuning guidelines, we refer to Markovsky et al. (2023). Select-DPC adds onto these the design parameters $N_{\text{cols}}$ as well as the method of data selection. In the following section, we will investigate two different data-selection methods which subsample a specified number of trajectories from a large offline library of trajectories observed during “training.”
SQP-MPC
@@
The iterative refinement of the solution until convergence is inspired by the model-based SQP-MPC algorithm for nonlinear systems Diehl et al. (2002). We direct the interested reader to Appendix B for a detailed discussion of SQP-MPC and the relationship to Select-DPC. We note that Select-DPC’s implicit linearization is performed in trajectory space. In the model-based case, linearizing in trajectory space corresponds to a linear time-varying parametrization of the nonlinear dynamics for the MPC. This linear, time-varying parametrization contrasts with an LTI formulation in which the Jacobians are only evaluated at the current “state” $y(k)$ and kept fixed along the horizon.
Computational complexity
@@
Solving the convex quadratic optimization problem at the basis of Select-DPC involves repeatedly solving a symmetric system of linear equations Stellato et al. (2020). Hence, we expect the algorithm to scale with $\mathcal{O}(n^{3})$ where $n$ indicates the number of optimization variables which is directly proportional to the number of trajectories $N_{\text{cols}}$ selected from the full data set. Hence, from a computational complexity standpoint, we wish to perform the optimization with as few trajectories as possible. Additionally, the dominant computation in the data-selection is the sorting of the data according to the relevance to $\tilde{\tau}$ which scales with $\mathcal{O}(n_{d}\log n_{d})$, which we postulate is less significant than the computational complexity of the optimization problem. While Select-DPC solves a convex quadratic program and is smaller than full-data DPC or nonlinear DPC, it still requires more compute than Sampling-Based MPC methods. There exist DPC variants which reduce the number of optimization variables in the predictive control problem Favoreel & De Moor (1999); Breschi et al. (2023) but the comparison of these methods is outside the scope of this study. For an overview of existing methods we refer to Verheijen et al. (2023).